OpenStudy (anonymous):

Can someone explain how I would do this? Determine the value of y for the system of equations. x − 3y = 4 2x + y = 8

OpenStudy (cwrw238):

you can do this by elimination first multiply the first equation by - 1 then adding will eliminate x solve for y then plug this value of y back into the origintl equation ( first one will do) so you can solve for x

OpenStudy (texaschic101):

x - 3y = 4 --->(-2) 2x + y = 8 ----------- -2x + 3y = -8 (result of multiplying by -2) 2x + y = 8 -----------add 4y = 0 y = 0 now sub 0 in for y in either equation 2x + y = 8 2x + 0 = 8 2x = 8 x = 4 check.. x - 3y = 4 4 - 3(0) = 4 4 - 0 = 4 4 = 4 (correct) x = 4 and y = 0

OpenStudy (cwrw238):

* correction multiply by -2 (!!!!!)

OpenStudy (anonymous):

Thanks so much!! @texaschic101 @cwrw238

OpenStudy (cwrw238):

yw

OpenStudy (texaschic101):

do you have any questions at all ?

OpenStudy (anonymous):

No I understood what you wrote @texaschic101

OpenStudy (texaschic101):

good to hear :)